OpenStudy (anonymous):

A square field is to be hydro-seeded. The field has a diagonal length of 225 yards. How many square feet need to be hydro-seeded?

OpenStudy (campbell_st):

a square is a special rhombus, where all angles are right angles. The area of a rhombus  x and y are the diagonal lengths and the area of a rhombus is \[A = \frac{1}{2} \times x \times y\] in a square x = y using this information you can find the area of the square. but 1st change 225 yards to feet. hope it helps

OpenStudy (anonymous):

Thank you!